# QCD corrections of $e^{+}e^{-}\to J/\psi+c+\bar{c}$ using the principle of maximum conformality

### v1

@@ 3.3 Renormalization scale setting for the NLO total cross section
where $n_{f}$ represents the number of active flavors, which is related to the QCD $\beta_{0}$-function through $\beta_{0}=11-2/3n_{f}$. The perturbative coefficients can be divided into conformal and non-conformal terms Mojaza:2012mf ; Brodsky:2013vpa , and then the NLO total cross section (18) transforms to
@@
where the conformal coefficients $r_{1,0}=c_{1}$, $r_{2,0}=c_{2,0}+33c_{2,1}/2$ and the non-conformal coefficient $r_{2,1}=-3c_{2,1}/4$.
The PMC provides a systematic and process-independent manner to fix conventional scheme-and-scale ambiguities. It is a kind of resummation, e.g. it resums all the known type of $\beta$-terms with the help of RGE and determines the correct magnitude of $\alpha_{s}$, whose precision is determined by the given order of the perturbative series. As for the present case, after applying the PMC, the NLO-level non-conformal $\beta_{0}$-terms will be resummed to achieve an effective strong coupling $\alpha_{s}(Q_{*})$, in which $Q_{*}$ is usually called as the PMC scale and it can be determined at the leading-log (LL) accuracy
@@
4 Numerical results and discussions
@@
$m_{c}=1.5\pm 0.1\,{\rm GeV},\;\sqrt{s}=10.6\,{\rm GeV},\alpha=1/137,|R_{J/\psi}(0)|^{2}=0.918\,{\rm GeV}^{3},$ (23)
where the electromagnetic coupling constant $\alpha$ is taken from Ref. Gong:2009ng . The input value for $|R_{J/\psi}(0)|^{2}$ is obtained by using $|R_{J/\psi}(0)|^{2}=\Gamma(J/\psi\to e^{+}e^{-})9m^{2}_{J/\psi}/[16\alpha^{2}(1-4C_{F}\alpha_{s}/\pi)]$, where $m_{J/\psi}=2m_{c}$ and $\Gamma(J/\psi\to e^{+}e^{-})=5.53~{}{\rm KeV}$ ParticleDataGroup:2022pth . For other values of $m_{c}$, the value of $|R_{J/\psi}(0)|^{2}$ should be multiplied by $(m_{c}/1.5~{}{\rm GeV})^{2}$. We use the package RunDec3 Herren:2017osy to evaluate the running QCD coupling constant $\alpha_{s}(\mu_{R})$ at two-loop accuracy. According to $\alpha_{s}(M_{Z})=0.1179$ ParticleDataGroup:2022pth , we obtain $\alpha_{s}(2m_{c})=0.252$.
4.1 Total cross section
@@
$\sigma_{\rm NLO}=c_{1}(\mu_{R})a_{s}^{2}(\mu_{R})+\left[c_{2}(2m_{c})+2c_{1}\beta_{0}\ln\frac{\mu_{R}^{2}}{4m_{c}^{2}}\right]a_{s}^{3}(\mu_{R}),$ (24)
where $c_{1}=391.589$ and $c_{2}(2m_{c})=16231-580.686n_{f}$. Starting from the scale-dependent initial series (24), one conventionally sets $\mu_{R}=2m_{c}$ to achieve its central value and varies $\mu_{R}\in[m_{c},4m_{c}]$ to estimate its scale uncertainty. For convenience, we call the results under this choice as conventional ones. After applying the PMC, following the above procedures (18-21), we obtain the following scale invariant series:
@@
Table 1: Contributions of the $\alpha_{s}^{2}$-terms and the $\alpha_{s}^{3}$-terms to the NLO total cross sections (in unit: pb) of $e^{+}e^{-}\to J/\psi+c+\bar{c}$ under conventional (Conv.) and PMC scale-setting approaches, respectively. The PMC predictions are scale invariant. For conventional results, their central values are for $\mu_{R}=2m_{c}$ and their errors are for $\mu_{R}\in[m_{c},4m_{c}]$.
We show how the LO and NLO total cross sections change with the renormalization scale $\mu_{R}$ in Fig. 4  [^1] Here as a cross-check of our predictions for the initial NLO series, we have found that if taking the same values for the input parameters, we will obtain exactly the same conventional predictions of Refs. Zhang:2006ay ; Gong:2009ng .. The LO terms of conventional series are highly scale dependent. By including the NLO QCD corrections, the scale dependence of conventional series becomes smaller due to the cancellation of scale dependence among different orders, which is however still sizable, e.g. the net scale error is still $\sim 74\%$ for $\mu_{R}\in[m_{c},4m_{c}]$. As for the PMC prediction, it becomes scale independent by using the $\{\beta_{0}\}$-terms to set the correct magnitude of $\alpha_{s}$. We also give separate contributions of the $\alpha_{s}^{2}$-terms and the $\alpha_{s}^{3}$-terms to the NLO total cross sections under conventional (Conv.) and PMC scale-setting approaches in Table 1. Table 1 shows that the $\alpha_{s}^{3}$-terms contribute significantly in both two cases, emphasizing the importance of the NLO-terms. The relative importance of $\alpha_{s}^{2}$-terms and $\alpha_{s}^{3}$-terms is $1:71\%$ for $\mu_{R}=2m_{c}$ under conventional scale-setting approach, which changes to $1:43\%$ under PMC scale-setting approach. Thus the PMC series not only removes the scale dependence but also has a better pQCD convergence, showing the importance of using a proper scale-setting approach.
@@
The uncertainty caused by $\Delta\alpha_{s}(M_{Z})$ is obtained by using $\alpha_{s}(M_{Z})=0.1179\pm 0.0010$ ParticleDataGroup:2022pth . And we obtain
@@
Eqs. (28, 29) show that the PMC prediction is more sensitive to the value of $\Delta\alpha_{s}(M_{Z})$. This is reasonable since the PMC uses the RGE-involved terms of the process to set the effective value of $\alpha_{s}$, thus its resultant pQCD series is more sensitive to the $\alpha_{s}$ running behavior that is governed by the RGE and the preciseness of the reference value $\alpha_{s}(M_{Z})$.
It is also helpful to have an estimation of the contributions from the UHO terms. For the purpose, we adopt the Padé approximation approach (PAA) Basdevant:1972fe ; Samuel:1992qg ; Samuel:1995jc , which provides a practical way for promoting a finite series to an analytic function, to do such an estimation. Detailed procedures for such an estimation can be found in Ref.Du:2018dma . For the present case, the estimated $\alpha_{s}^{4}$-terms (e.g. the NNLO-terms) of the conventional series and PMC series are
@@
To compare with the experimental data, we should also include the QED contribution as well as the two-photon contribution via the process $e^{+}e^{-}\to 2\gamma^{*}\to J/\psi+c+\bar{c}$. Since the experimental measurements are for the prompt $J/\psi+c+\bar{c}$ production, we should also consider the feed-down contributions from the higher excited charmonium states such as $e^{+}e^{-}\to\psi(2S)+c+\bar{c}\to J/\psi+c+\bar{c}+X$ and $e^{+}e^{-}\to\chi_{cJ}+c+\bar{c}\to J/\psi+c+\bar{c}+X$. By using $\Gamma(\psi(2S)\to e^{+}e^{-})=5.53~{}{\rm KeV}$ and the branching ratio for the $\psi(2S)\to J/\psi X$ transition fraction $Br[\psi(2S)\to J/\psi X]=61.4\%$ ParticleDataGroup:2022pth , we obtain the contribution from $e^{+}e^{-}\to\psi(2S)+c+\bar{c}\to J/\psi+c+\bar{c}+X$ is $0.371~{}\sigma_{\rm NLO}$. The other contributions have been discussed in Ref. Zhang:2006ay , which lead to a shift of $0.069$ pb to the total cross section. We will give the comparison with experimental data in the following section.
4.2 Differential cross section
The differential distributions contain more information than the integrated cross sections. Experimentally, the $J/\psi$ three-momentum and angular distributions for $e^{+}e^{-}\to J/\psi+c+\bar{c}$ have been measured by the Belle collaboration Belle:2009bxr . It is interesting to see the differential distributions of the process $e^{+}e^{-}\to J/\psi+c+\bar{c}$.
@@ 4.2 Differential cross section
4.3 Self-consistency of PMC to integrated and differential cross sections
@@
We show the PMC NLO predictions of the total cross sections in Table 2, which are derived either from the differential series ($d\sigma/dP_{J/\psi}$, $d\sigma/d|\cos\theta|$ and $d\sigma/dz$) or from the series (25). The upper indexes ${P_{J/\psi}}$, ${|\cos\theta|}$ and ${z}$ denote the ones derived from integrating over the differential cross sections $d\sigma/dP_{J/\psi}$, $d\sigma/d|\cos\theta|$ and $d\sigma/dz$, respectively. Table 2 shows that all those PMC predictions are consistent with each other. This shows that the PMC predictions can also get the self-consistent results for the differential and total cross sections. But there are differences in the details of getting the same conclusion. There are slight differences, which are less than ±$\pm 1\%$, among the results obtained by integrating over the differential distributions with the result obtained from the total series (25). Such slight difference is reasonable, 1) the differences are due to the resummation of different types of $\{\beta_{i}\}$-terms, and are caused by the unknown $\{\beta_{i}\}$-types from the unknown NNLO and higher order terms; 2) The effects of the differences of the $\{\beta_{i}\}$-terms will result in the differences of the PMC scales, which generally suffer from both exponential and $\alpha_{s}$ suppressions. So the net value for such difference could be small. For the present case of $e^{+}e^{-}\to J/\psi+c+\bar{c}$, we observe that the differences are about ±$\pm 0.003$ pb, which are higher-order effects and are under well control  [^2] For some cases, certain large logarithmic terms may appear in boundary region of the phase space, which may spoil the convergent behavior of the perturbative series for some specific phase-space points. And special resummation, such as the threshold resummation Yan:2023mjj , need to be done so as to achieve a more reliable PMC prediction.. And for the present case, their magnitudes are much smaller than the predicted NNLO magnitude ±$\pm 0.047$ pb. Thus we can conclude that the PMC can be applied to any pQCD calculable observable via a self-consistent way of perturbative theory, and it can achieve more precise fixed-order perturbative predictions that are free of renormalization scheme-and-scale ambiguities.
5 Summary
@@
Figure 11: The prompt total cross sections of $e^{+}e^{-}\to J/\psi+c+\bar{c}$ under conventional and PMC scale-setting approaches, respectively. The Belle measured value Belle:2009bxr is given as a comparison.
In Fig. 11, we present a comparison of the prompt total cross section of $e^{+}e^{-}\to J/\psi+c+\bar{c}$ with the Belle measured value, $\sigma^{\rm Exp.}(e^{+}e^{-}\to J/\psi+c+\bar{c})=0.74^{+0.113}_{-0.120}$ pb, whose errors are squared averages of the systematic and statistical errors given in Eq.(2). Fig. 11 shows that a better agreement with the data can be achieved by applying the PMC scale-setting approach.
After applying the PMC scale-setting approach, one can determine the correct magnitude of the effective $\alpha_{s}$, resulting in a more convergent pQCD series that is free of divergent renormalon terms. Thus, a reliable, self-consistent and precise pQCD prediction can be achieved. We have also done a first consistency test of the PMC approach to deal with the total and differential observables. As a byproduct, the typical momentum flow of the process may be small in specific phase space, the consistency property of the PMC production makes its differential series be a useful platform for testing the correctness of the low-energy $\alpha_{s}$ models.
Acknowledgments: We thank Stanley J. Brodsky for helpful discussions. This work was supported in part by the National Natural Science Foundation of China under Grant Nos. 12175025, 12347101, 12247129, 11975242, and 12135013; by the Graduate Research and Innovation Foundation of Chongqing, China under Grant No. ydstd1912; by the Foundation of Chongqing Normal University under Grant No. 24XLB015.

### v2

@@ III.3 The NLO cross sections and the PMC
where $n_{f}$ represents the number of active flavors, which is related to the QCD $\beta_{0}$-function through $\beta_{0}=11-2/3n_{f}$. The perturbative coefficients can be divided into conformal and non-conformal terms [46, 47], and then the NLO total cross section (18) transforms to
@@
where the LO and NLO conformal coefficients $r_{1,0}=c_{1}$ and $r_{2,0}=c_{2,0}+33c_{2,1}/2$, and the NLO non-conformal coefficient $r_{2,1}=-3c_{2,1}/4$.
The PMC provides a process-independent manner to fix conventional scheme-and-scale ambiguities. In practice, it is a kind of resummation, e.g. it resums all the known type of $\beta$-terms with the help of RGE and determines the correct magnitude of $\alpha_{s}$, whose precision is determined by the given order of the perturbative series. As for the present case, following the standard PMC scale-setting procedures, the NLO-level non-conformal $\beta_{0}$-terms will be resummed to achieve an effective strong coupling $a_{s}(Q_{*})$, i.e.,
$a^{2}_{s}(Q_{*})$$\leftarrow$ $a^{2}_{s}(\mu_{R})\left[1+2\beta_{0}\frac{r_{i+1,1}}{r_{i,0}}a_{s}(\mu_{R})+\cdots\right],$ (21)
which is derived by using the RGE. The scale $Q_{*}$ is usually called as the PMC scale and it can be determined at the leading-log (LL) accuracy
@@
IV Numerical results and discussions
@@
$m_{c}=1.5\pm 0.1\,{\rm GeV},\;\sqrt{s}=10.6\,{\rm GeV},$
$\alpha=1/137,|R_{J/\psi}(0)|^{2}=0.918\,{\rm GeV}^{3},$ (24)
where the electromagnetic coupling constant $\alpha$ is taken from Ref. [35]. The input value for $|R_{J/\psi}(0)|^{2}$ is obtained by using $|R_{J/\psi}(0)|^{2}=\Gamma(J/\psi\to e^{+}e^{-})9m^{2}_{J/\psi}/[16\alpha^{2}(1-4C_{F}\alpha_{s}/\pi)]$, where $m_{J/\psi}=2m_{c}$ and $\Gamma(J/\psi\to e^{+}e^{-})=5.53~{}{\rm KeV}$ [73]. For other values of $m_{c}$, the value of $|R_{J/\psi}(0)|^{2}$ should be multiplied by $(m_{c}/1.5~{}{\rm GeV})^{2}$. We use the package RunDec3 [74] to evaluate the running of the QCD coupling constant $\alpha_{s}(\mu_{R})$ at two-loop accuracy. According to $\alpha_{s}(M_{Z})=0.1179$ [73], we obtain $\alpha_{s}(2m_{c})=0.252$.
IV.1 Total cross section
@@
$\sigma_{\rm NLO}=c_{1}a_{s}^{2}(\mu_{R})+\bigg{[}c_{2}(2m_{c})+2c_{1}\beta_{0}\ln\frac{\mu_{R}^{2}}{4m_{c}^{2}}\bigg{]}a_{s}^{3}(\mu_{R}),$ (25)
where $c_{1}=391.589$ and $c_{2}(2m_{c})=16231-580.686n_{f}$. Starting from the scale-dependent initial series (25), one conventionally sets $\mu_{R}=2m_{c}$ to achieve its central value and varies $\mu_{R}\in[m_{c},4m_{c}]$ to estimate its scale uncertainty. For convenience, we call the results under this choice as conventional ones. After applying the PMC, following the above procedures (18-22), we obtain the following scale invariant series:
@@
Table 1: Contributions of the $\alpha_{s}^{2}$-terms and the $\alpha_{s}^{3}$-terms to the NLO total cross sections (in unit: pb) of $e^{+}e^{-}\to J/\psi+c+\bar{c}$ under conventional (Conv.) and PMC scale-setting approaches, respectively. The PMC predictions are scale invariant. The conventional results are scale dependent, and their central values are for $\mu_{R}=2m_{c}$ and their errors are for $\mu_{R}\in[m_{c},4m_{c}]$.
We show how the LO and NLO total cross sections change with the renormalization scale $\mu_{R}$ in Fig. 4  [^1] Here as a cross-check of our predictions for the initial NLO series, we have found that if taking the same values for the input parameters, we will obtain exactly the same conventional predictions of Refs. [34, 35].. The LO terms of conventional series are highly scale dependent. By including the NLO QCD corrections, the scale dependence of conventional series becomes smaller due to the cancellation of scale dependence among different orders, which is however still sizable, e.g. the net scale error is still $\sim 74\%$ for $\mu_{R}\in[m_{c},4m_{c}]$. As for the PMC prediction, it becomes scale independent by using the $\{\beta_{0}\}$-terms to set the correct magnitude of $\alpha_{s}$. We also give separate contributions of the $\alpha_{s}^{2}$-terms and the $\alpha_{s}^{3}$-terms to the NLO total cross sections under conventional (Conv.) and PMC scale-setting approaches in Table 1. Table 1 shows that the $\alpha_{s}^{3}$-terms contribute significantly in both two cases, emphasizing the importance of the NLO-terms. The relative importance of $\alpha_{s}^{2}$-terms and $\alpha_{s}^{3}$-terms is $1:71\%$ for $\mu_{R}=2m_{c}$ under conventional scale-setting approach, which changes to $1:43\%$ under PMC scale-setting approach. Thus the PMC series not only removes the scale dependence but also has a better pQCD convergence, showing the importance of using a proper scale-setting approach.
@@
The uncertainty caused by $\Delta\alpha_{s}(M_{Z})$ is obtained by using $\alpha_{s}(M_{Z})=0.1179\pm 0.0010$ [73]. And we obtain
@@
Eqs. (29, 30) show that the PMC prediction is more sensitive to the value of $\Delta\alpha_{s}(M_{Z})$. This is reasonable since the PMC uses the RGE-involved terms of the process to set the effective value of $\alpha_{s}$, thus its resultant pQCD series is more sensitive to the $\alpha_{s}$ running behavior that is governed by the RGE and the preciseness of the reference value $\alpha_{s}(M_{Z})$.
Since the NLO-terms are sizable, it is helpful to have an estimation of the contributions from the UHO terms. For the purpose, we adopt the Padé approximation approach (PAA) [75, 76, 77], which provides a practical way for promoting a finite series to an analytic function, to do such an estimation. Detailed procedures for such an estimation can be found in Ref.[78]. For the present case, the estimated $\alpha_{s}^{4}$-terms (e.g. the NNLO-terms) of the conventional series and PMC series are
@@
To compare with the experimental data, we should also include the QED contribution as well as the two-photon contribution via the process $e^{+}e^{-}\to 2\gamma^{*}\to J/\psi+c+\bar{c}$. Since the experimental measurements are for the prompt $J/\psi+c+\bar{c}$ production, we should also consider the feed-down contributions from the higher excited charmonium states such as $e^{+}e^{-}\to\psi(2S)+c+\bar{c}\to J/\psi+c+\bar{c}+X$ and $e^{+}e^{-}\to\chi_{cJ}+c+\bar{c}\to J/\psi+c+\bar{c}+X$. By using $\Gamma(\psi(2S)\to e^{+}e^{-})=5.53~{}{\rm KeV}$ and the branching ratio for the $\psi(2S)\to J/\psi X$ transition fraction $Br[\psi(2S)\to J/\psi X]=61.4\%$ [73], we obtain the contribution from $e^{+}e^{-}\to\psi(2S)+c+\bar{c}\to J/\psi+c+\bar{c}+X$ is $0.371~{}\sigma_{\rm NLO}$. The other contributions have been discussed in Ref. [34], which lead to a shift of $0.069$ pb to the total cross section. We will give the comparison with experimental data in the following section.
IV.2 Differential cross section
The differential distributions contain more information than the integrated cross sections. Experimentally, the $J/\psi$ three-momentum and angular distributions for $e^{+}e^{-}\to J/\psi+c+\bar{c}$ have been measured by the Belle collaboration [15]. It is interesting to see the differential distributions of the process $e^{+}e^{-}\to J/\psi+c+\bar{c}$.
@@ IV.2 Differential cross section
IV.3 Self-consistency of PMC to integrated and differential cross sections
@@
We show the PMC NLO predictions of the total cross sections in Table 2, which are derived either from the differential series ($d\sigma/dP_{J/\psi}$, $d\sigma/d|\cos\theta|$ and $d\sigma/dz$) or from the series (26). The upper indexes ${P_{J/\psi}}$, ${|\cos\theta|}$ and ${z}$ denote the ones derived from integrating over the differential cross sections $d\sigma/dP_{J/\psi}$, $d\sigma/d|\cos\theta|$ and $d\sigma/dz$, respectively. Table 2 shows that all those PMC predictions are consistent with each other. This shows that the PMC predictions can also get the self-consistent results for the differential and total cross sections. But there are differences in the details of getting the same conclusion. There are slight differences, which are less than ±$\pm 1\%$, among the results obtained by integrating over the differential distributions with the result obtained from the total series (26). Such slight difference is reasonable, 1) the differences are due to the resummation of different types of $\{\beta_{i}\}$-terms, and are caused by the unknown $\{\beta_{i}\}$-types from the unknown NNLO and higher order terms; 2) The effects of the differences of the $\{\beta_{i}\}$-terms will result in the differences of the PMC scales, which generally suffer from both exponential and $\alpha_{s}$ suppressions. So the net value for such difference could be small. For the present case of $e^{+}e^{-}\to J/\psi+c+\bar{c}$, we observe that the differences are about ±$\pm 0.003$ pb, which are higher-order effects and are under well control  [^2] For some cases, certain large logarithmic terms may appear in boundary region of the phase space, which may spoil the convergent behavior of the perturbative series for some specific phase-space points. And special resummation, such as the threshold resummation [79], need to be done so as to achieve a more reliable PMC prediction.. And for the present case, their magnitudes are much smaller than the predicted NNLO magnitude ±$\pm 0.047$ pb. Thus we can conclude that the PMC can be applied to any pQCD calculable observable via a self-consistent way of perturbative theory, and it can achieve more precise fixed-order perturbative predictions that are free of renormalization scheme-and-scale ambiguities.
V Summary
@@
Figure 11: The prompt total cross sections of $e^{+}e^{-}\to J/\psi+c+\bar{c}$ under conventional and PMC scale-setting approaches, respectively. The Belle measured value [15] is given as a comparison.
In Fig. 11, we present a comparison of the prompt total cross section of $e^{+}e^{-}\to J/\psi+c+\bar{c}$ with the Belle measured value, $\sigma^{\rm Exp.}(e^{+}e^{-}\to J/\psi+c+\bar{c})=0.74^{+0.113}_{-0.120}$ pb, whose errors are squared averages of the systematic and statistical errors given in Eq.(2). Fig. 11 shows that a better agreement with the experimental data can be achieved by applying the PMC.
After applying the PMC scale-setting approach, one can determine the correct magnitude of the effective $\alpha_{s}$, generally resulting in a more convergent pQCD series that is free of divergent renormalon terms. Thus, a reliable, self-consistent and precise pQCD prediction can be achieved. We have also done a first consistency test of the PMC approach to deal with the total and differential observables. As a byproduct, the typical momentum flow of the process may be small in specific phase space, the consistency property of the PMC production makes its differential series be a useful platform for testing the correctness of the low-energy $\alpha_{s}$ models.
Acknowledgments: We thank Stanley J. Brodsky for helpful discussions. This work was supported in part by the National Natural Science Foundation of China under Grant Nos. 12175025, 12347101, 12247129, 11975242, and 12135013; by the Graduate Research and Innovation Foundation of Chongqing, China under Grant No. ydstd1912; by the Foundation of Chongqing Normal University under Grant No. 24XLB015; by the Fundamental Research Funds for the Central Universities under Grant No. 2024CDJXY022.